A 30-year maturity bond making annual coupon payments with a coupon rate of 14.3% has duration of 11.34 years and convexity of 185.7. The bond currently sells at a yield to maturity of 8%.

a. Find the price of the bond if its yield to maturity falls to 7%. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

b. What price would be predicted by the duration rule? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

c. What price would be predicted by the duration-with-convexity rule? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

d-1. What is the percent error for each rule? (Enter your answer as a positive value. Do not round intermediate calculations. Round "Duration Rule" to 2 decimal places and "Duration-with-Convexity Rule" to 3 decimal places.)


d-2. What do you conclude about the accuracy of the two rules?

  • The duration rule provides more accurate approximations to the actual change in price.

  • The duration-with-convexity rule provides more accurate approximations to the actual change in price.


e-1. Find the price of the bond if it's yield to maturity rises to 9%. (Do not round intermediate calculations. Round your answer to 2 decimal places.)


e-2. What price would be predicted by the duration rule? (Do not round intermediate calculations. Round your answer to 2 decimal places.)


e-3. What price would be predicted by the duration-with-convexity rule? (Do not round intermediate calculations. Round your answer to 2 decimal places.)


e-4. What is the percent error for each rule? (Do not round intermediate calculations. Round "Duration Rule" to 2 decimal places and "Duration-with-Convexity Rule" to 3 decimal places.)


e-5. Are your conclusions about the accuracy of the two rules consistent with parts (a) – (d)?

  • Yes

  • No
Answer #1
Bond Price =∑ [(Annual Coupon)/(1 + YTM)^k]     +   Par value/(1 + YTM)^N
                   k=1
                  K =30
Bond Price =∑ [(14.3*1000/100)/(1 + 8/100)^k]     +   1000/(1 + 8/100)^30
                   k=1
Bond Price = 1709.24
a)
New bond price @ YTM =7
                  K = N
Bond Price =∑ [(Annual Coupon)/(1 + YTM)^k]     +   Par value/(1 + YTM)^N
                   k=1
                  K =30
Bond Price =∑ [(14.3*1000/100)/(1 + 7/100)^k]     +   1000/(1 + 7/100)^30
                   k=1
Bond Price = 1905.86

b

b)
New bond price @ YTM =7 using duration
Mod.duration prediction = -Mod. Duration*Yield_Change*Bond_Price
=-11.34*-0.01*1709.24
=193.827816
New bond price = bond price+Modified duration prediction
=1709.24+193.827816
=1903.07

c

New bond price @ YTM =7 using duration and convexity
Convexity adjustment = 0.5*convexity*Yield_Change^2*Bond_Price
=0.5*185.7*-0.01^2*1709.24
=15.8702934
New bond price = bond price+Mod.duration pred.+convex. Adj.
=1709.24+193.83+-15.87
=1918.94

d-1

Percentage error for YTM =7 and duration rule
Difference in price predicted and actual
=predicted price-actual price
=1903.067816-1905.86
=-2.792
Please ask remaining parts seperately